OpenStudy (anonymous):

if $1200 is deposited in an account that pays 6.5% interest, what is the difference in the ammount after 4 yearsbetween the ammount earned if the princial is compounded annualy and the amount earned calculated using simple interest

OpenStudy (tkhunny):

Have you considered calculating the two values?

OpenStudy (anonymous):

yes I have tried

OpenStudy (tkhunny):

Demonstrate. I = Prt ==> I = 1200*0.065*4 = 312.00 ==> 1200 + 312 = 1512 Okay, there's the Simple Interest. Show the Annual Compounding.

OpenStudy (tkhunny):

A = P(1 + r)^t ==> What?

OpenStudy (anonymous):

that's not my problem ( and i already know how to do the formula) my problem is how to find out if 1.12 or if 0.72 or 2.54 or 3.76 is my answer.

OpenStudy (tkhunny):

Demonstrate the use of the formula.

OpenStudy (tkhunny):

It is important for you to use the formula in public, so you can more effectively refute those four silly answer choices.

OpenStudy (tkhunny):

They are silly answer choices. We just need to prove it. Did you get them from a worksheet or online bunch of choices?

OpenStudy (anonymous):

how are they silly answer choices? it is a question I have to solve but I cant figure out the answer, don't have to be rude about it

OpenStudy (tkhunny):

I'm never rude. Honesty and directness are often mistaken for rudeness. They are silly answer choices because they are not anywhere near the correct answer based on the problem statement. All I have asked you to do is demonstrate the usage of the formula. Nothing rude about that. Once you do that, we can find the correct answer and be on our way.

OpenStudy (anonymous):

well then i cant find the right answer because those are the only 4 choices I have for answers

OpenStudy (tkhunny):

You'll have to take that up with the author of the question. If you demonstrate the use of the formula, we can send you on your way with the right answer - noting that it is NOT one of the four given choices.

OpenStudy (tkhunny):

Oh, how I wish I could get you to show some of YOUR work. Too bad. You could learn more.

OpenStudy (mathstudent55):

The formula for simple interest is: I = Prt where I = interest earned P = principal (beginning) amount r = interest rate written as a decimal t = number of years Can you use the simple interest formula, show all the numbers you use, and find the simple interest earned on $1200 at 6.5% for 4 years?

OpenStudy (mathstudent55):

Write the formula I = Prt On the next line, replace P, r, and t with the given values. Then find I. Can you do that and show it?

OpenStudy (mathstudent55):

I'm trying to help you. We can find an answer. Are you still interested?

OpenStudy (mathstudent55):

Are you serious?

OpenStudy (tkhunny):

Showing Your Work - Looks like you are trying to learn. Not Showing Your Work - Look like you are trying to cheat - whether you are trying to cheat or not. It's an appearance thing.

OpenStudy (anonymous):

ok I already did the question by myself

OpenStudy (tkhunny):

Yup. That's exactly what cheaters say. Showing your work is the ONLY way to DEMONSTRATE otherwise. I have seen NONE of YOUR work on ANY of the questions you have posted. You need to do better than that.

OpenStudy (mathstudent55):

@Mavckenzie Out of curiosity, what answer did you get?